Check if the following utility functions represent the same preferences: u = x+y, v = x3 +

y3, w = -1/(x+y). Give reasons for your answer.


1
Expert's answer
2021-10-29T11:13:38-0400

Using marginal rate of substitution approach which says that if the different utility function gives us the Same rate of substitution when they represent the same preferences

"U=X+Y, V=X^3+Y^3, W=-\frac{1}{X+Y}"

"U=X+Y\\


MUX=\frac{\delta U}{\delta x}=1\\

MUY=\frac{\delta U}{\delta Y}=1\\


MRS=\frac{-MUX}{MUY}\\MRS=-1....(1)"


"V=X^3+Y^3\\MUX=3X^2\\MUY=3Y^2"

"MRS=\frac{-MUX}{MUY}\\MRS=\frac{-X^2}{Y^2}\\MRS=\frac-(\frac{X}{Y})^2....(2)"

Now for

"W=-1(X+Y)^{-1}\\
MUX=\frac{\delta W}{\delta X}=(X+Y)^{-2}\\MUY=\frac{\delta W}{\delta Y}(X+Y)^{-2}\\

MRS=\frac{-MUX}{MUY}\\

MRS=\frac{-(X+Y)^{-2}}{(X+Y)^{-2}}\\MRS=-1...(3)"


Equition 1 and 3 show that U and W represent the same preferences but V does not.